# Differentiable Grouped Feedback Delay Networks for Learning Coupled Volume Acoustics

### v1

@@ VII Conclusion
We showed that the RIRs generated by DiffGFDN are analogous to those produced by the CS model, which is a state-of-the-art method. Compared to the CS model, which relies on modal reverberators and equalisers, DiffGFDN is more computationally efficient and generalises to unseen positions. We proposed a subband architecture with multiple parallel GFDNs, and evaluated it against the CS model on a simulated three-room dataset. Although the network produced slightly higher EDC errors than the CS model, it achieved lower EDR errors. Lower EDC errors were achieved when more RIRs were used for training. However, training with only 3030% of RIRs in the dataset produced less than $0.6$ dB difference in EDC error compared to the best performing models in Table I in all but the $4$ kHz band; in the $4$ kHz band, the worst EDC error difference compared to the best performing model was $1.3$ dB.
Future work will explore optimised implementations using decimating filterbanks, introducing coupling in the lower frequencies, spatial sampling requirements near room transitions, extensions to directional late reverberation modelling, and formal perceptual evaluations in real measured environments.

### v2

@@ VIII Conclusion
We showed that the RIRs generated by DiffGFDN are analogous to those produced by the CS model, which is a state-of-the-art method. Compared to the CS model, which relies on modal reverberators and equalisers, DiffGFDN is more computationally efficient and interpolates to unseen positions. We proposed a subband architecture with multiple parallel decoupled GFDNs, and evaluated it in an ablation study against other configurations, and against the CS model on a simulated three-room dataset. We also studied the effect of varying the amount of training data on the octave-band EDC errors in a fixed hold-out test set.
Future work will explore generalisation to unseen geometries, using decimating filterbanks, spatial sampling requirements near room transitions and formal perceptual evaluations in real measured environments.
IX Acknowledgments
The authors wish to thank Georg Götz for the helpful and insightful discussions around the common slope model.
The Aalto University School of Electrical Engineering funded the work of the second author and supported it with the HUCE infrastructure. The computational resources were provided by the Aalto Science-IT project.